# QFNN-FFD: Quantum Federated Neural Network for Financial Fraud Detection

### v2

@@ Abstract
This study introduces the Quantum Federated Neural Network for Financial Fraud Detection (QFNN-FFD), a cutting-edge framework merging Quantum Machine Learning (QML) and quantum computing with Federated Learning (FL) for financial fraud detection. Using quantum technologies’ computational power and the robust data privacy protections offered by FL, QFNN-FFD emerges as a secure and efficient method for identifying fraudulent transactions within the financial sector. Implementing a dual-phase training model across distributed clients enhances data integrity and enables superior performance metrics, achieving precision rates consistently above 95%. Additionally, QFNN-FFD demonstrates exceptional resilience by maintaining an impressive 80% accuracy, highlighting its robustness and readiness for real-world applications. This combination of high performance, security, and robustness against noise positions QFNN-FFD as a transformative advancement in financial technology solutions and establishes it as a new benchmark for privacy-focused fraud detection systems. This framework facilitates the broader adoption of secure, quantum-enhanced financial services and inspires future innovations that could use QML to tackle complex challenges in other areas requiring high confidentiality and accuracy.
Index Terms:
Quantum Neural Network, Quantum Federated Learning, Quantum Machine Learning, Fraud Detection, Finance
@@ I Introduction
In the rapidly evolving financial technology landscape, privacy is a fundamental pillar, crucial for upholding the trust and integrity of financial transactions and services [1]. As digital transactions become more prevalent, the volume of sensitive data handled by financial institutions grows exponentially, making robust privacy measures indispensable [2]. The emergence of Quantum Machine Learning (QML) marks a transformative era [3, 4, 5], promising unprecedented computational capabilities by exploiting quantum physics [6], while simultaneously raising pivotal concerns about privacy and data security. This paper introduces the Quantum Federated Neural Network for Financial Fraud Detection (QFNN-FFD), a pioneering framework that influences the quantum-enhanced processing power of Quantum Computing (QC) with the privacy-preserving attributes of Federated Learning (FL). The synergy of QML with FL jointly improves the efficiency and accuracy of detecting fraudulent activities, while safeguarding sensitive financial data against the ever-looming threats of breaches and unauthorized access.
QFNN-FFD represents a significant leap forward in applying quantum technologies to real-world economic challenges and sets a new benchmark for privacy-centric approaches in the fintech domain. By deploying this framework, financial institutions can potentially harness the unique advantages of QC—such as rapid processing of large datasets—while also benefiting from the decentralized nature of FL, which keeps sensitive data localized and reduces the risk of central points of failure. As shown in Fig. 1, Quantum Federated Learning (QFL) has shown superior performance in various fields [7, 8, 9], prompting our decision to implement it in finance. Our framework has demonstrated its capability to enhance both accuracy and privacy protection through comparative analysis with existing models works [10, 11, 12]. This approach meets and often surpasses current industry standards, providing a scalable, secure framework that adapts seamlessly to diverse operational environments while maintaining high accuracy in fraud detection under various conditions.
Refer to caption (a)(b)
Figure 1: Comparison of ML and FL accuracies in classical and QC contexts across various fields and experiments. Panel (a)illustrates the performance of different experiments within the finance sector. Panel (b) compares QML with QFL across four domains: healthcare, IoT, computer v, and finance. In classical computing contexts, FL generally demonstrates superior performance compared to ML [13, 14]. In QC contexts, QFL exhibits slight improvements over QML [15, 16, 17]. These findings highlight the potential of QFL and provide a compelling rationale for its adoption, particularly in the finance sector.
Our contributions significantly impact the fintech sector by providing a scalable, secure framework that adapts to various operational environments while maintaining high accuracy in fraud detection under different conditions.
@@
Our novel contributions, encapsulating a comprehensive workflow that significantly enhances fintech security measures, are listed as follows and shown in Fig. 2:
@@ II Background and Related Works
The progression of FL into the QC domain has precipitated the inception of QFL [17, 21]. This methodology exploits quantum mechanics’ distinctive properties to augment privacy and computational efficiency. In [22], the study delineated the first fully operational QFL framework capable of processing exclusively quantum data. This innovation indicated the establishment of the inaugural quantum federated dataset, facilitating the collaborative learning of quantum circuit parameters by quantum clients in a decentralized manner—a cornerstone in adapting quantum technologies to federated contexts.
Subsequently, the notion of dynamic QFL frameworks was advanced in [8], which introduced the Slimmable QFL. This framework was designed to adapt to varying network conditions and constraints on computing resources by dynamically modulating the training parameters of Quantum Neural Networks (QNNs). Empirical studies demonstrated that SlimQFL sustained superior classification accuracy under fluctuating conditions compared to conventional static QFL methods. Moreover, integrating secure protocols, such as blind quantum computation, into distributed learning environments enhanced the confidentiality aspect of QFL.
The research outlined in [23] proposed a quantum protocol that leveraged the computational capacities of remote quantum servers while safeguarding the privacy of the underlying data. This protocol proved its robustness against prevalent security threats, including gradient attacks, rendering it especially beneficial in domains that demand boosted security for distributed learning tasks.
@@
Advancing into financial fraud, significant research has been conducted to apply QC and QML in detecting financial fraud. In [12], they developed quantum protocols for anomaly detection, applying them to credit card fraud. They compared quantum kernel methods to classical ML benchmarks, revealing the potential for quantum approaches to achieve superior precision significantly as the number of qubits increased. This demonstrated that quantum methods can become increasingly advantageous with the scaling of quantum systems.
Furthermore, in [25], they explored using a Quantum Support Vector Machine (QSVM) for real-world financial data, presenting one of the first end-to-end applications of QML in the financial sector. Their findings highlighted the complementary nature of QSVM to classical techniques, offering new methods for feature exploration that enhanced fraud detection accuracy.
As the application of QML in fraud detection advances, several innovative approaches have emerged. For instance, in [26], they explored using QML models, including the Variational Quantum Classifier (VQC) and different QNNs. These models showed promising results in classifying fraud and non-fraud transactions, demonstrating QML’s potential in financial applications. Despite their success, these models faced challenges such as needing more efficient quantum algorithms and the ability to handle larger and more complex datasets.
In [27], the study addressed the latency in traditional fraud detection systems by implementing a QML approach using a Support Vector Machine (SVM) enhanced with quantum annealing solvers. This approach significantly outperformed traditional models’ speed and accuracy, particularly on time-series data from bank loans, which are typically highly imbalanced.
In [28], they discussed a hybrid model that combines QNNs with classical neural networks to enhance fraud detection capabilities. This study implemented two distinct methods: a hybrid quantum-classical neural network and topological data analysis for noise reduction and improved classification accuracy. Such hybrid models leverage the computational power of quantum devices while maintaining the versatility and robustness of classical ML frameworks.
Generative adversarial networks (GANs) have also been adapted to quantum settings to tackle the instability and inefficiency of classical sampling methods. [29] introduced variational quantum-classical Wasserstein GANs (WGANs), which incorporated a hybrid quantum-classical generator with a classical discriminator. This model was effective on a credit card fraud dataset, providing competitive performance with classical counterparts in terms of F1 score.
Further advancing the field, in [30], they presented an approach using data re-uploading techniques to train single-qubit classifiers that perform comparably to classical models under similar training conditions. This study highlights the potential for QML to achieve significant results with minimal quantum resources, opening new avenues for efficient quantum computations.
Moreover, in [31] and [32], they highlighted the real-time challenges in fraud detection. They utilized quantum annealing to develop frameworks that enhance the speed of fraud detection, addressing the critical issue of timely response in fraudulent transaction identification.
@@ III-A QNN Circuit Design and QFL Integration
Figure 4: General schematic of a QML model workflow. The process begins with qubits in the zero state $(|0\rangle)$. The qubits undergo data encoding to represent the input data in quantum form. Then, a parametrized quantum circuit, $U(\theta)$, transforms the qubit states, where $\theta$ represents tunable parameters. The transformed quantum states are measured, converting quantum information into classical output. This output is evaluated using a predefined loss function, and a classical optimization algorithm iteratively adjusts $\theta$ to minimize the loss, thereby refining the QML model’s performance.
The QFNN-FFD framework operates on data distributed across $N$ clients, each possessing a subset of the overall dataset, thereby preventing the need for central data aggregation and enhancing data privacy. Training of the QFNN-FFD is directed in a federated manner, where local models on each client are independently trained using their data subsets.
In the local model, the first step is to encode classical data into quantum states through angle encoding. Each data feature $x_{i,j}$ from the vector $\mathbf{x}_{i}$ of client $i$ is mapped onto two rotation angles, $\theta_{i,j}$ for the $R_{y}$ rotation and $\phi_{i,j}$ for the $R_{z}$ rotation. These rotations are then applied to the qubits sequentially to modify both their phase and orientation:
@@ III-C Parameter Aggregation and Model Evaluation
Following local optimization, each client’s parameters $\theta_{i}$ are transmitted to a central server. They are aggregated through a simple averaging process to update the global model parameters $\theta$. This cyclic process of local optimization and global aggregation iteratively enhances the QFNN-FFD’s performance, evaluated on a global validation set for generalizability and efficacy. The mathematical foundation of parameter optimization within the QFNN-FFD employs the Adam optimizer, adjusting $\theta$ as
@@
The globally updated parameters are redistributed to all clients for further training, cycling through local optimization and global aggregation to progressively improve the QFNN-FFD’s performance. This iterative process enhances computational efficiency and maintains strict privacy standards.
Integrating QML with FL in our QFNN-FFD framework fosters the high-efficiency processing of complex financial data and upholds stringent data privacy standards. This dual advantage, coupled with the model’s mathematical rigor and strategic parameter optimization, positions the QFNN-FFD as an effective tool in the fight against financial fraud, marking a significant leap forward in applying QC to real-world challenges in the financial sector.
@@ IV-A Experimental Setup
Our QFNN-FFD is implemented using PennyLane for model architecture and Qiskit for simulating quantum noise [38, 39], enabling a realistic QC environment. The framework, structured around four qubits, employs the Adam optimizer ($\eta$=0.1) across dual training phases—local and global—with up to 100 iterations for each across 15 clients. This setup is characterized by 32 initially random parameters, which are optimized through evaluations on a training set comprising 115,386 instances (80% of the total dataset of 144,233 instances) and a validation set comprising 28,847 instances, which is 20% of the total dataset. We focus on binary classification accuracy and MSE as key metrics. Operational deployment occurs within an environment characterized by a configuration consisting of 4 virtual CPUs (vCPUs), 25 gigabytes (GB) of RAM, and a single NVIDIA Tesla V100 virtual GPU (vGPU). This setup offers a balance of processing power, memory capacity, and advanced GPU acceleration—crucial factors for efficiently handling the intensive computations required by our QFNN-FFD framework.
@@ IV-B Accuracy and Loss Analysis
Refer to caption123Stable plateau indicatinghigh performance3Rapid improvement phase2Initial spike inloss reduction
@@
IV-C Quantum Noise Analysis
@@
IV-D Comparison with Existing Works
TABLE I: Comparison of QML frameworks on different financial fraud datasets.
Reference Precision Recall F1-Score Accuracy
[10] 84 84.44 75.68 83.92
[11] 96.1 79.5 86 94.5
[12] 90
Our QFNN-FFD 95 96 95 95
@@ IV-D Comparison with Existing Works
IV-E Discussion
Our results show that the framework achieves high validation accuracy, maintains low loss across various operational conditions, and exhibits resilience against diverse quantum noise models. Such robustness underlines the framework’s suitability for real-world QC environments known for their integral noise issues. In direct comparison with existing quantum and classical models, QFNN-FFD surpasses typical performance metrics, making it a superior choice for fraud detection. This performance is particularly notable given the framework’s integration of privacy-preserving FL, which safeguards sensitive financial data during detection. This dual benefit of enhanced accuracy and increased data privacy sets QFNN-FFD apart as a leading solution for secure and effective fraud detection in the fintech industry. Furthermore, the framework’s ability to maintain high performance under various noise conditions suggests its potential for broader applications beyond financial services, including sectors where data sensitivity and security are paramount. Integrating advanced quantum computational capabilities with robust privacy features positions QFNN-FFD as a scalable solution for future challenges in secure data processing and analysis.

### v3

@@ Abstract
This work introduces a Quantum Federated Neural Network for Financial Fraud Detection (QFNN-FFD), an advanced framework merging Quantum Machine Learning (QML) and quantum computing with Federated Learning (FL) for financial fraud detection. Using quantum technologies’ computational power and the robust data privacy protections offered by FL, QFNN-FFD emerges as a secure and efficient method for identifying fraudulent transactions within the financial sector. Implementing a dual-phase training model across distributed clients enhances data integrity and enables superior performance metrics, achieving precision rates consistently above 95%. Additionally, QFNN-FFD demonstrates exceptional resilience by maintaining an impressive 80% accuracy, highlighting its robustness and readiness for real-world applications. This combination of high performance, security, and robustness against noise positions QFNN-FFD as a transformative advancement in financial technology solutions and establishes it as a new approach for privacy-focused fraud detection systems. This framework facilitates the broader adoption of secure, quantum-enhanced financial services and inspires future innovations that could use QML to tackle complex challenges in other areas requiring high confidentiality and accuracy.
@@ I Introduction
In the rapidly evolving financial technology landscape, privacy is a fundamental pillar, crucial for upholding the trust and integrity of financial transactions and services [1]. As digital transactions become more prevalent, the volume of sensitive data handled by financial institutions grows exponentially, making robust privacy measures indispensable [2]. The emergence of Quantum Machine Learning (QML) marks a transformative era [3, 4, 5, 6, 7, 8, 9], promising computational capabilities by exploiting quantum physics [10], while simultaneously raising pivotal concerns about privacy and data security. This paper introduces the Quantum Federated Neural Network for Financial Fraud Detection (QFNN-FFD), a framework that integrates the quantum-enhanced processing power of Quantum Computing (QC) with the privacy-preserving attributes of Federated Learning (FL). The synergy of QML with FL jointly improves the efficiency and accuracy of detecting fraudulent activities, while safeguarding sensitive financial data against the ever-looming threats of breaches and unauthorized access.
QFNN-FFD demonstrates the potential of quantum technologies in addressing real-world economic challenges and sets a new benchmark for privacy-centric approaches in the fintech domain. By deploying this framework, financial institutions can potentially employ the advantages of QC—such as the potential rapid processing of large datasets—while also benefiting from the decentralized nature of FL, which keeps sensitive data localized and reduces the risk of central points of failure. As shown in Fig. 1, Quantum Federated Learning (QFL) has shown superior performance in various fields [11, 12, 13, 14], prompting our decision to implement it in finance. Our framework has demonstrated its capability to enhance both accuracy and privacy protection through comparative analysis with existing models [15, 16, 17]. This approach meets and often surpasses current industry standards, providing a scalable, secure framework that adapts seamlessly to diverse operational environments while maintaining high accuracy in fraud detection under various conditions.
Figure 1: Comparison of ML and FL accuracies in classical and QC contexts across various fields and experiments. Panel (a) illustrates the performance of different experiments within the finance sector. Panel (b) compares QML with QFL across four domains: healthcare, IoT, computer vision, and finance. In classical computing contexts, FL generally demonstrates superior performance compared to ML [18, 19]. In QC contexts, QFL exhibits slight improvements over QML [20, 21, 22]. These findings highlight the potential of QFL and provide a compelling rationale for its adoption, particularly in the finance sector.
@@
Our contributions significantly impact the fintech sector by providing a secure framework that adapts to various operational environments while maintaining high accuracy in fraud detection under different conditions, and can be listed as follows and shown in Fig. 2:
@@ II Background and Related Works
The progression of FL into the QC domain has precipitated the inception of QFL [22, 26, 27]. This methodology exploits quantum mechanics’ distinctive properties to augment privacy and computational efficiency. In [28], the study delineated the first fully operational QFL framework capable of processing exclusively quantum data. This innovation indicated the establishment of the inaugural quantum federated dataset, facilitating the collaborative learning of quantum circuit parameters by quantum clients in a decentralized manner—a cornerstone in adapting quantum technologies to federated contexts.
Subsequently, the notion of dynamic QFL frameworks was advanced in [12], which introduced the Slimmable QFL. This framework was designed to adapt to varying network conditions and constraints on computing resources by dynamically modulating the training parameters of Quantum Neural Networks (QNNs). The research outlined in [29] proposed a quantum protocol that used the computational capacities of remote quantum servers while safeguarding the privacy of the underlying data.
@@
Advancing into financial fraud, significant research has been conducted to apply QC and QML in detecting financial fraud. In [17], they developed quantum protocols for anomaly detection, applying them to credit card fraud. Furthermore, in [31], they explored using a Quantum Support Vector Machine (QSVM) for real-world financial data, presenting one of the first end-to-end applications of QML in the financial sector. As the application of QML in fraud detection advances, several innovative approaches have emerged. For instance, in [32], they explored using QML models, including the Variational Quantum Classifier (VQC) and different QNNs. These models showed promising results in classifying fraud and non-fraud transactions, demonstrating QML’s potential in financial applications. In [33], the study addressed the latency in traditional fraud detection systems by implementing a QML approach using a Support Vector Machine (SVM) enhanced with quantum annealing solvers. In [34], they discussed a hybrid model that combines QNNs with classical neural networks to enhance fraud detection capabilities.
Generative adversarial networks (GANs) have also been adapted to quantum settings to tackle the instability and inefficiency of classical sampling methods. [35] introduced variational quantum-classical Wasserstein GANs (WGANs), which incorporated a hybrid quantum-classical generator with a classical discriminator. This model was effective on a credit card fraud dataset, providing competitive performance with classical counterparts in terms of F1 score. Further advancing the field, in [36], they presented an approach using data re-uploading techniques to train single-qubit classifiers that perform comparably to classical models under similar training conditions. Moreover, in [37] and [38], they highlighted the real-time challenges in fraud detection.
@@ III-A QNN Circuit Design and QFL Integration
Figure 5: General schematic of a QML model workflow. The process begins with qubits in the zero state $(|0\rangle)$. The qubits undergo data encoding to represent the input data in quantum states. Then, a parametrized quantum circuit, $U(\theta)$, transforms the qubit states, where $\theta$ represents tunable parameters. The transformed quantum states are measured, converting quantum information into classical output. This output is evaluated using a predefined loss function, and a classical optimization algorithm iteratively adjusts $\theta$ to minimize the loss, thereby refining the QML model’s performance.
The QFNN-FFD framework operates on data distributed across $N$ clients, with each client possessing an Identically and Independently Distributed (IID) subset of the overall dataset. This uniform distribution ensures that all clients train the models under similar data conditions, which prevents the need for central data aggregation and enhances data privacy.
Training of the QFNN-FFD is directed in a federated manner, where local models on each client are independently trained using their data subsets.
In the local model, the first step is to encode classical data into quantum states through angle encoding. Each data feature $x_{i,j}$ from the vector $\mathbf{x}_{i}$ for client $i$ is mapped onto two rotation angles, $\theta_{i,j}$ for the $R_{y}$ rotation and $\phi_{i,j}$ for the $R_{z}$ rotation. These rotations are then applied to the qubits sequentially to modify both their phase and orientation:
@@ III-B Optimization and Training Process
For gradient computation in the QNN circuit, we employ the parameter-shift rule [43]. This method is particularly well-suited for QML models because it provides an exact estimation of the gradient, avoiding the issues associated with finite differences and stochastic gradient methods. The gradient of the QNN circuit parameter $\theta_{i}$ is given by:
$\frac{\partial L}{\partial\theta_{i}}=\frac{L(\theta_{i}+\frac{\pi}{2})-L(\theta_{i}-\frac{\pi}{2})}{2},$ (7)
where $L(\theta)$ denotes the loss function evaluated at the parameter-shifted values.
@@ III-C Parameter Aggregation and Model Evaluation
Following local optimization, each client’s parameters $\theta_{i}$ are transmitted to a central server. They are aggregated through a simple averaging process to update the global model parameters $\theta$. This cyclic process of local optimization and global aggregation iteratively enhances the QFNN-FFD’s performance, which is evaluated on a global validation set for generalizability and efficacy. The mathematical foundation of parameter optimization within the QFNN-FFD employs the Adam optimizer, adjusting $\theta$ as
@@
The globally updated parameters are redistributed to all clients for further training, cycling through local optimization, and global aggregation to progressively improve the QFNN-FFD’s performance. This iterative process enhances computational efficiency and maintains strict privacy standards.
Integrating QML with FL in our QFNN-FFD framework fosters the high-efficiency processing of complex financial data and upholds stringent data privacy standards. This dual advantage, coupled with the model’s mathematical rigor and strategic parameter optimization, positions the QFNN-FFD as an effective tool in the fight against financial fraud, marking substantial progress in applying QC to real-world challenges in the financial sector.
III-D Computational Complexity Analysis
The computational complexity of our framework arises from both the depth of the QNN circuit and the FL process. Our QNN implementation processes each data point through multiple layers of single-qubit rotations and entanglement operations. This results in a per-sample computational complexity of $\mathcal{O}(nL)$, where $n$ is the number of qubits and $L$ is the total number of layers. Gradient computation using the parameter-shift rule requires two additional circuit evaluations per parameter, leading to a per-iteration complexity of $\mathcal{O}(P)$, where $P$ is the total number of parameters and is proportional to $nL$. Considering $T$ local iterations over a dataset of size $D$ per client, the overall computational complexity per client becomes $\mathcal{O}(TDP)$. The federated averaging process adds minimal overhead, with a communication complexity of $\mathcal{O}(P)$ per client per round. Despite the intensive quantum computations, the workload is effectively distributed across clients, and efficient gradient computation ensures the framework remains computationally feasible for practical applications in FFD.
@@ IV-A Experimental Setup
Our QFNN-FFD is implemented using PennyLane for model architecture and Qiskit for simulating quantum noise [45], enabling a realistic QC environment. The framework, structured around four qubits, employs the Adam optimizer ($\eta$=0.1) across dual training phases—local and global—with up to 100 iterations for each across 15 clients. This setup is characterized by 32 initially random parameters, which are optimized through evaluations on a training set comprising 115,386 instances (80% of the total dataset of 144,233 instances) and a validation set comprising 28,847 instances, which is 20% of the total dataset. We focus on binary classification accuracy and MSE as key metrics. Operational deployment occurs within an environment characterized by a configuration consisting of 4 virtual CPUs (vCPUs), 25 gigabytes (GB) of RAM, and a single NVIDIA Tesla V100 virtual GPU (vGPU). This setup offers a balance of processing power, memory capacity, and advanced GPU acceleration—crucial factors for efficiently handling the intensive computations required by our QFNN-FFD framework.
@@ IV-B Accuracy and Loss Analysis
IV-C Local Client Validation Accuracy Analysis
The validation accuracy trends across individual clients reveal important aspects of our framework’s performance on distributed datasets (batches). As shown in Fig. 8, the initial variability in accuracies suggests that each client’s model interacts differently with its local data, influenced by the unique characteristics of the data and the stochastic elements in the training process, such as random initialization and the adaptive optimizer’s behavior. As training progresses, the results show a clear convergence pattern. By the 70th iteration, the validation accuracies of most clients stabilize, with the majority achieving accuracy levels around or above 0.9. This indicates that our framework successfully synchronizes the learning trajectories across the distributed clients, leading to a consistent and robust overall performance. The increasing mean accuracy, as shown by the trend line, underscores our framework’s ability to harmonize client models, even in a decentralized setting. Despite the overall positive trend, minor fluctuations in accuracy persist for some clients, particularly in the later stages of training. These fluctuations likely result from the heterogeneous nature of the data across clients or the sensitivity of individual models to specific features. While these variations do not detract significantly from the overall convergence, they highlight the necessity of continuous monitoring and potential fine-tuning to ensure that all clients achieve optimal performance.
Figure 8: Validation accuracy trends for individual clients across selected iterations (10, 30, 50, 70, 100). Each bar represents the accuracy of a specific client at a given iteration, with the dashed black line indicating the mean accuracy trend across all clients. The figure illustrates the initial variability in client performance, followed by a convergence toward higher accuracy levels as training progresses, with most clients stabilizing around or above 0.9 accuracy by the 100th iteration. Minor fluctuations observed in some clients highlight the influence of data heterogeneity and model sensitivity on individual learning outcomes.
IV-D Quantum Noise Analysis
@@
IV-E Comparison with Existing Works
TABLE I: Comparison of QML frameworks on financial fraud datasets.
Reference Precision Recall F1-Score Accuracy Privacy [15] 84 84.44 75.68 83.92 $\times$ [16] 96.1 79.5 86 94.5 $\times$ [17] 90 $\times$ Our QFNN-FFD 95 96 95 95
The datasets used in the compared studies have been carefully selected using the same attributes as those employed to train our QFNN-FFD, transaction volume, diversity in the variation of values for transactions and merchant categories, and the incidence rates of fraud cases. These are the main selection criteria that ensure uniformity in the level of complexity of datasets and applicability to the particular challenges of financial fraud detection. This careful selection ensures that our comparative analysis is appropriately contextualized and reflects real-world transactional environments and their complexities.
@@ IV-E Comparison with Existing Works
IV-F Discussion
Our results show that the framework achieves high validation accuracy, maintains low loss across various operational conditions, effectively harmonizes distributed learning tasks across clients, and exhibits resilience against diverse quantum noise models. Such robustness underlines the framework’s suitability for real-world QC environments known for their integral noise issues. In direct comparison with existing quantum and classical models, QFNN-FFD surpasses typical performance metrics, making it a superior choice for fraud detection. This performance is particularly notable given the framework’s integration of privacy-preserving FL, which safeguards sensitive financial data during detection. This dual benefit of enhanced accuracy and increased data privacy sets QFNN-FFD apart as a leading solution for secure and effective fraud detection in the fintech industry. Furthermore, the framework’s ability to maintain high performance under various noise conditions suggests its potential for broader applications beyond financial services, including sectors where data sensitivity and security are paramount. Integrating advanced quantum computational capabilities with robust privacy features positions QFNN-FFD as a scalable solution for future challenges in secure data processing and analysis.